# Displacement memory for flyby

### v4

@@ Abstract
Zel’dovich and Polnarev, in their seminal paper [1] suggested that a gravitational wave generated by flyby would merely displace the particles. We confirm their prediction numerically by fine-tuning the wave profile proposed by Gibbons and Hawking [2], and then analytically for its approximation by a Pöschl-Teller potential. Higher-order derivative profiles proposed for gravitational collapse, etc [2] are shortly discussed.
Annals of Physics (accepted)
@@ I Introduction
The Displacement Memory Effect (DM) for gravitational waves generated by flyby was proposed by Zel’dovich and Polnarev ZelPol (see also GibbHaw71 ; BraTho ; BraGri ; Sachs62 ; Christo ), extending previous work on the Velocity Memory Effect (VM) Ehlers ; Sou73 ; GriPol . Our initial investigations ShortMemory ; LongMemory ; EZHRev hinted at, though, that for randomly chosen parameters particles initially at rest would fly apart with non-zero velocity, as illustrated, e.g. in FIG. # 12. of ref. EZHRev . However closer recent scrutiny DM-1 ; Jibril ; Benin indicates that for certain “magical” values of the wave parameters reminiscent of quantum numbers we do get DM, highlighted by the vanishing of the velocity outside the Wavezone,
@@
This paper extends these results to a more physical realm with particular attention at the analytically solvable Pöschl-Teller -version DM-1 ; Benin ; PTeller ; Chakra . Then we generalize the flyby result to gravitational collapse and other similar cases by studying models with higher-order derivatives of the (Gaussian or Pöschl-Teller profile) proposed in BraTho ; BraGri ; GibbHaw71 .
@@ II Flyby: transverse motion
We first study what happens when the profile is the derivative of a Gaussian, proposed by Gibbons and Hawking for flyby in GibbHaw71 ,
@@
Figure 1: The null geodesics for the first derivative of the Gaussian in (II.1) shown in the background exhibit, for a randomly choosen fixed amplitude $k$ but with different initial conditions, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(0.5,0.5,0.5)},\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(1,1,1)},\,{\color[rgb]{0,0.35,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.35,0}(1.5,1.5,1.5)}$, the velocity effect (VM).
We start with the massless case, $\mathfrak{m}=0$. For randomly chosen amplitude the null geodesics seem to show VM. We study first the transverse motion; that in the $V$-direction will be considered in sect.VI. Numerical fine-tuning then indicates that, unexpectedly, for certain “magical” critical values $k=k_{crit}$ we do get (approximate) DM for both transverse coordinates, as illustrated in FIGs.2 and 3. In fact, we found a well separated discrete series of critical amplitudes, $k_{m}$, each of which corresponding to a unique DM wave labeled by an integer $m$,
@@
Our plots suggest that for each such critical amplitude $k_{crit}$ the Wavezone contains a unique DM trajectory composed approximately of an integer number of half-waves, consistently with our observations in DM-1 . The novelty is that DM is now obtained for both coordinates. This is unlike as we had for the simple Gaussian we studied shown in FIG.20 of ref. DM-1 for which DM could be found, by fine-tuning, in the attractive, but not in the repulsive sector that we call therefore “half-DM”. This is behavior is not accidental, as we shall see later.
@@
Figure 3: DM trajectories with (approximate) wave number ${\bf m=1}$ for the flyby profile (II.1), presented separately for $X^{1}$ and for $X^{2}$. The outgoing values depend strongly on the initial conditions, choosen here as ${\color[rgb]{0,0.35,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.35,0}(0.1,5)},\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(0.2,10)},\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(0.3,15)}$. The parameters were chosen to emphasise the reciprocity relation (IV.20), to be discussed in sec.IV.
@@
The behavior for the derivative-Gaussian profile (II.1) for $m=1$ and with different initial conditions is shown in FIG.3 for comparison with FIG.1. In the $U\leq 0$ region the force is repulsive, and in $U\geq 0$ it is attractive. The strong dependence of the trajectory on the initial position we observe on FIG.3 follows from the “reciprocity constraint” (IV.20), as it will be discussed in section IV.
For generic wave parameters, the particle absorbs some energy from the passing wave Bondi57 ; Ehlers ; exactsol . DM implies in turn zero cumulated energy of the solution in FIG.2, as confirmed in FIG.5. The energy absorption of the underlying NR model Eisenhart ; DBKP ; DGH91 for $k_{crit}$ both in the low and high amplitude regimes, is shown in FIG.6. We shall return to the relation with the underlying NR system in sec.VI.
@@ III Pöschl-Teller model for flyby: numerical study
allowed us to find analytic solutions by solving the Pöschl-Teller -modified Sturm-Liouville equation DM-1 ; Jibril ; Chakra ; Benin
@@
Our clue is that the DM conditions (I.1) [or $X(U=\mp\infty)\to\mathop{\rm const.}\nolimits$] require that the solution extends to $t=\mp 1$, — yielding a quantization condition (see the Conclusion). The parameter $m$ must be a natural number,
@@
where $m$ is an a priori real number. Thus (III.8) is a Legendre eqn. with a linear perturbation. For $t\to\pm 1$ both equations become Legendre equations but with different coefficients of the linear term, see FIGs. 8 and 9.
@@ IV.1 Results
$\begin{split}k_{1}=29&.056007,\\ k_{2}=84&.097092,\\ k_{3}=166&.608873,\end{split}\qquad\begin{split}k_{4}&=276.61564,\\ k_{5}&=414.12126,\\ k_{6}&=579.12680\,.\end{split}$ (IV.16)
@@ IV.2 Asymptotic values
$X^{1}(\infty)=-83.541230\,.$ (IV.18)
For $k=\pm k_{2}$ the analogous results are
$X^{1}(\infty)=2176.4880,\qquad X^{2}(\infty)=0.00045945\,.$ (IV.19)
The necessity to keep so many digits follows from the reciprocity constraint,
$X^{1}(\infty)X^{2}(\infty)=1\,,$ (IV.20)
which comes itself from $U$-reversal symmetry ,
$U\leftrightarrow-U\,,$ (IV.21)
which implies in turn $X^{2}(\infty)=X^{1}(-\infty)/X^{1}(\infty)=1/X^{1}(\infty)$ with our normalization.
@@ IV.3 The $k_{m}\leftrightarrow m$ relation
The following, purely numerically obtained formulas,
$k_{m}=b_{0}w_{m}+b_{1}+b_{2}/w_{m}\,,{\quad\text{\small where}\quad}w_{m}=m(m+1),$ (IV.22)
and
$b_{0}=13.750365,\qquad b_{1}=1.61433036,\qquad b_{2}=-0.11798667\,$ (IV.23)
work with rather good precision.
@@ IV.4 Coefficient of the $\ln(1-z)$ singularity
$\sum_{n=k}^{\infty}\frac{z^{n}}{n^{2}(n-1)(n-2)\dots(n-k+1)}=-c_{k}\ln(1-z)+r_{k}.$ (IV.25)
@@
$c_{1}=-\zeta-\frac{\zeta^{2}}{2}-\frac{\zeta^{3}}{3}+\dots\qquad r_{1}=\quad\frac{\pi^{2}}{6}-\zeta-\frac{\zeta^{2}}{4}-\frac{\zeta^{3}}{9}+\dots\quad$ (IV.26)
$c_{2}=\frac{\zeta^{2}}{2}+\frac{\zeta^{3}}{3}+\dots\qquad\qquad\;r_{2}=2-\frac{\pi^{2}}{6}-\zeta+\frac{\zeta^{2}}{4}+\frac{\zeta^{3}}{9}+\dots$ (IV.27)
@@
$\sum_{k=0}^{\infty}A_{k}c_{k}(\zeta)=A_{0}\sum_{n=0}^{\infty}u_{m}\zeta^{m}.$ (IV.29)
@@
$u_{0}=1\qquad u_{1}=p\qquad u_{2}=\frac{p^{2}}{4}\qquad u_{3}=-\frac{p^{2}}{18}+\frac{p^{3}}{36}\,.$ (IV.30)
@@
$s_{0}=1\qquad s_{1}=-p\qquad s_{2}=\frac{p^{2}}{4}\qquad s_{3}=-\frac{p^{2}}{18}-\frac{p^{3}}{36}\,.$ (IV.31)
We see that they agree up to a sign change. This is not accidental. First of all we see that the differential operator of the confluent Heun equation ${\cal D}X=0$ is
${\cal D}=z(z-1)\frac{{\rm d}^{2}}{{\rm d}z^{2}}+(2z-1)\frac{{\rm d}}{{\rm d}z}+p(2z-1),$ (IV.32)
@@
${\cal D}=\zeta(\zeta-1)\frac{{\rm d}^{2}}{{\rm d}\zeta^{2}}+(2\zeta-1)\frac{{\rm d}}{{\rm d}\zeta}-p(2\zeta-1)\,,$ (IV.33)
i.e., it is the same operator with $p\to-p$. Let us write the Ansatz
$X=-c(\zeta)\ln\zeta+r(\zeta),$ (IV.34)
where we assume that the coefficient functions $c(\zeta)$ and $r(\zeta)$ are analytic at $\zeta=0$ (power series). Applying the differential operator to this Ansatz we get
${\cal D}c=0{\quad\text{\small and}\quad}{\cal D}r=2(\zeta-1)\frac{{\rm d}c}{{\rm d}\zeta}+c\,.$ (IV.35)
From (IV.35) we then see that $c(\zeta)$ satisfies the $(-p)$ Heun equation and we find that
$c(\zeta)=A_{0}{\rm HeunC}\big{[}-p,-2p,1,1,0,\zeta\big{]}.$ (IV.36)
@@ IV.5 Half-waves
Of course, perfect half-waves are sinusoidal. For $X(U)=\sin U$ the nodes are at
$U_{1}=0,\qquad U_{2}=\pi,\qquad U_{3}=2\pi\dots$ (IV.37)
@@
${\rm between\;}U_{1}\ {\rm and\ }U_{2}:\quad X(U)>0,\qquad X^{\prime\prime}(U)=-\sin U<0\,.$ (IV.38)
@@
${\rm between\;}U_{2}\ {\rm and\ }U_{3}:\quad X(U)<0,\qquad X^{\prime\prime}(U)=-\sin U>0\,.$ (IV.39)
@@ IV.5.1 Pöschl-Teller
$\frac{{\rm d}^{2}}{{\rm d}U^{2}}X(U)=-\frac{m(m+1)}{\cosh^{2}U}X(U)$ (IV.40)
@@
$U_{r}={\rm arctanh}\,y_{r}\,,\qquad r=1,2,\dots m\,.$ (IV.41)
With our normalization $X(-\infty)=1$ and so for $U<U_{1}$ $X(U)>0$ :
$U_{1}<U<U_{2}:\qquad X(U)<0,\quad X^{\prime\prime}(U)>0\,.$ (IV.42)
@@
$U_{2}<U<U_{3}:\qquad X(U)>0,\quad X^{\prime\prime}(U)<0\,.$ (IV.43)
The solution is concave, bends back downwards until $U_{3}$. This alternating pattern repeats itself until $U_{m}$. The solution in each of the intervals is monotonic until the (in this interval unique) local minimum/maximum is reached; then it is again monotonic towards the next node. This looks like a distorted half-wave .
The main reason of the half-wave -behaviour is that $X(U)$ and $X^{\prime\prime}(U)$ have opposite sign.
Between $U_{1}$ and $U_{m}$ there are exactly $m-1$ such distorted half-waves . It is easy to see that this $m-1$ counting is correct by plotting
$m=1\qquad X(U)=-P_{1}(\tanh U)=-\tanh U\,.$ (IV.44)
This solution is monotonic, there is obviously no half-wave as seen in FIG.#7 of DM-1 . Similarly one can plot
$m=2\qquad X(U)=-\frac{1}{2}+\frac{3}{2}\tanh^{2}U.$ (IV.45)
There is a single half-wave between $U_{1}=-{\rm arctanh}(1/\sqrt{3})$ and $U_{2}={\rm arctanh}(1/\sqrt{3})$, as seen in FIG.#8 of DM-1  .
@@ IV.5.2 Derived Pöschl-Teller
$\frac{{\rm d}^{2}}{{\rm d}U^{2}}X(U)=-k\tanh U\frac{1}{\cosh^{2}U}X(U)\,.$ (IV.46)
As we have seen, for the half-wave -behaviour it is crucial that $X$ and $X^{\prime\prime}$ have opposite sign. In this case this is satisfied only for $U>0$. However, it turns out that this is not a problem. In our normalization $U(-\infty)=1$ and so the solution starts off positive. But since initially $X^{\prime\prime}$ is also positive, the solution curve is convex, so it bends upwards and does not vanish for negative $U$. After crossing $U=0$ the curve becomes concave and now bends downwards until it reaches the first node $U_{1}$. All nodes of $X^{1}$ are positive. This means that all our qualitative considerations for the PT case are valid also here: there are exactly $(m-1)$ half-waves between $U_{1}$ and $U_{m}$.

### v6

@@ Abstract
Zel’dovich and Polnarev, in their seminal paper [1] extending previous work on the Memory Effect, suggested that a gravitational wave generated by flyby would merely displace the particles. We confirm their prediction numerically by fine-tuning the wave profile proposed by Gibbons and Hawking [2], and then analytically for its approximation by a Pöschl-Teller potential. Higher-order derivative profiles proposed for gravitational collapse, etc [2] are shortly discussed.
Annals Phys. 473 (2025), 169890. doi:10.1016/j.aop.2024.169890 [arXiv:2407.10787 [gr-qc]]
@@ I Introduction
The Displacement Memory Effect (DM) for gravitational waves generated by flyby was proposed by Zel’dovich and Polnarev ZelPol (see also GibbHaw71 ; BraTho ; BraGri ; Sachs62 ; Christo ), extending previous work on the Velocity Memory Effect (VM) Ehlers ; Sou73 ; GriPol . Our previous investigations ShortMemory ; LongMemory ; EZHRev hinted, though, at that for randomly chosen parameters particles initially at rest would fly apart with non-zero velocity, as illustrated, e.g. in FIG. # 12. of ref. EZHRev . However closer recent scrutiny DM-1 ; Jibril ; ZZHflyby indicates that for certain “magical” values of the wave parameters reminiscent of quantum numbers we do get DM, highlighted by the vanishing of the velocity outside the Wavezone,
@@
This paper extends these results to a more physical realm with particular attention at the analytically solvable Pöschl-Teller -version DM-1 ; ZZHflyby ; PTeller ; Chakra . Then we generalize the flyby result to gravitational collapse, etc by studying models with higher-order derivatives of the (Gaussian or Pöschl-Teller profile) proposed in BraTho ; BraGri ; GibbHaw71 .
@@ II Flyby: transverse motion
II.1 Flyby profile: derivative of Gaussian
The profile for flyby proposed by Gibbons and Hawking in GibbHaw71 is,
@@ II.1 Flyby profile: derivative of Gaussian
Figure 1: The null geodesics for the first derivative of the Gaussian in (II.1) shown in the background exhibit, for randomly choosen amplitude $k$, the velocity effect (VM).
We start with the massless case, $\mathfrak{m}=0$. At the first sight, the null geodesics show VM as seen in FIG.1. We study first the transverse motion; that in the $V$-direction will be considered in sect.VI. Numerical fine-tuning then indicates, unexpectedly, that for certain “magical” critical values $k=k_{crit}$ we do get (approximate) DM for both transverse coordinates, as illustrated in FIG.2. In fact we found a well separated discrete series of critical amplitudes, $k_{m}$, each of which corresponding to a unique DM wave labeled by an integer $m$,
@@
Our plots suggest that for each such critical amplitude $k_{crit}$ the Wavezone contains a unique DM trajectory composed approximately of an integer number of half-waves, consistently with our observations in DM-1 . The novelty is that DM is now obtained for both coordinates, whereas for the simple Gaussian, shown in FIG.20 of DM-1 , DM could be found in the attractive, but not in the repulsive sector; in DM-1 we called this “half-DM”.
@@
For generic wave parameters, the particle absorbs some energy from the passing wave Bondi57 ; Ehlers ; exactsol . DM implies in turn zero cumulated energy of the solution in FIG.2, as confirmed in FIG.4.
@@
It is instructive to record the energy absorption of the underlying NR model Eisenhart ; DBKP ; DGH91 for $k_{crit}$ both in the low and high amplitude regimes, shown in FIG.5.
@@ III Pöschl-Teller model for flyby: numerical study
allowed us to find analytic solutions by solving the Pöschl-Teller -modified Sturm-Liouville equation DM-1 ; Jibril ; Chakra ; ZZHflyby
@@
Our clue is that the DM conditions (I.1) [or $X(U=\mp\infty)\to\mathop{\rm const.}\nolimits$] require that the solution extends to $t=\mp 1$, — yielding a quantization condition: the parameter $m$ must be a natural number,
@@
where $m$ is an a priori real number. Thus (III.8) is a Legendre eqn. with a linear perturbation. For $t\to\pm 1$ both equations become Legendre equations but with different coefficients of the linear term, see FIGs. 7 and 8.
@@ IV.1 Results
$\begin{split}k_{1}=29&.0560068187,\\ k_{2}=84&.0970923581,\\ k_{3}=166&.6088726851,\end{split}\qquad\begin{split}k_{4}&=276.6156438862,\\ k_{5}&=414.1212560177,\\ k_{6}&=579.1267981111.\end{split}$ (IV.16)
@@ IV.2 Asymptotic values
$X^{1}(\infty)=-83.5412300512\,.$ (IV.18)
Of course, $X^{1}(\infty)X^{2}(\infty)=1$, due to the $U\leftrightarrow-U$ symmetry of the problem and our normalization. For $k=\pm k_{2}$ the analogous results are
[…]
@@ IV.3 The $k_{m}\leftrightarrow m$ relation
The following, purely numerically obtained formula,
$k_{m}=b_{0}w_{m}+b_{1}+b_{2}/w_{m}\,,$ (IV.20)
where $w_{m}=m(m+1)$ and
$b_{0}=13.750365230964,\qquad b_{1}=1.614330366353,\qquad b_{2}=-0.117986675402.$ (IV.21)
works with rather good precision.
@@ IV.4 Coefficient of the $\ln(1-z)$ singularity
$\sum_{n=k}^{\infty}\frac{z^{n}}{n^{2}(n-1)(n-2)\dots(n-k+1)}=-c_{k}\ln(1-z)+r_{k}.$ (IV.23)
@@
$c_{1}=-\zeta-\frac{\zeta^{2}}{2}-\frac{\zeta^{3}}{3}+\dots\qquad r_{1}=\quad\frac{\pi^{2}}{6}-\zeta-\frac{\zeta^{2}}{4}-\frac{\zeta^{3}}{9}+\dots\quad$ (IV.24)
$c_{2}=\frac{\zeta^{2}}{2}+\frac{\zeta^{3}}{3}+\dots\qquad\qquad\;r_{2}=2-\frac{\pi^{2}}{6}-\zeta+\frac{\zeta^{2}}{4}+\frac{\zeta^{3}}{9}+\dots$ (IV.25)
@@
$\sum_{k=0}^{\infty}A_{k}c_{k}(\zeta)=A_{0}\sum_{n=0}^{\infty}u_{m}\zeta^{m}.$ (IV.27)
@@
$u_{0}=1\qquad u_{1}=p\qquad u_{2}=\frac{p^{2}}{4}\qquad u_{3}=-\frac{p^{2}}{18}+\frac{p^{3}}{36}\,.$ (IV.28)
@@
$s_{0}=1\qquad s_{1}=-p\qquad s_{2}=\frac{p^{2}}{4}\qquad s_{3}=-\frac{p^{2}}{18}-\frac{p^{3}}{36}\,.$ (IV.29)
We see that they agree up to a sign change. This is not accidental. First of all we can see that the original differential operator of the confluent Heun equation ${\cal D}X=0$ is
${\cal D}=z(z-1)\frac{{\rm d}^{2}}{{\rm d}z^{2}}+(2z-1)\frac{{\rm d}}{{\rm d}z}+p(2z-1),$ (IV.30)
@@
${\cal D}=\zeta(\zeta-1)\frac{{\rm d}^{2}}{{\rm d}\zeta^{2}}+(2\zeta-1)\frac{{\rm d}}{{\rm d}\zeta}-p(2\zeta-1)$ (IV.31)
i.e. it is the same operator with $p\to-p$. Let us write the ansatz
$X=-c(\zeta)\ln\zeta+r(\zeta),$ (IV.32)
where we assume that the coefficient funcions $c(\zeta)$ and $r(\zeta)$ are analytic at $\zeta=0$ (power series). Applying the differential operator to this ansatz we get
${\cal D}c=0$ (IV.33)
and
${\cal D}r=2(\zeta-1)\frac{{\rm d}c}{{\rm d}\zeta}+c\,.$ (IV.34)
From (IV.33) we then see that $c(\zeta)$ satisfies the $-p$ Heun equation and we find that
$c(\zeta)=A_{0}{\rm HeunC}\big{[}-p,-2p,1,1,0,\zeta\big{]}.$ (IV.35)
@@ IV.5 Half-waves
Of course, the perfect half-waves (HW) are sinusoidal. For $X(U)=\sin U$ the nodes are at
$U_{1}=0,\qquad U_{2}=\pi,\qquad U_{3}=2\pi\dots$ (IV.36)
@@
${\rm between\;}U_{1}\ {\rm and\ }U_{2}:\quad X(U)>0,\qquad X^{\prime\prime}(U)=-\sin U<0\,.$ (IV.37)
@@
${\rm between\;}U_{2}\ {\rm and\ }U_{3}:\quad X(U)<0,\qquad X^{\prime\prime}(U)=-\sin U>0\,.$ (IV.38)
@@ IV.5.1 Pöschl-Teller
$\frac{{\rm d}^{2}}{{\rm d}U^{2}}X(U)=-\frac{m(m+1)}{\cosh^{2}U}X(U)$ (IV.39)
@@
$U_{r}={\rm arctanh}\,y_{r},\qquad r=1,2,\dots m\,.$ (IV.40)
With our normalization $X(-\infty)=1$ and so for $U<U_{1}$ $X(U)>0$, then:
$U_{1}<U<U_{2}:\qquad X(U)<0,\quad X^{\prime\prime}(U)>0\,.$ (IV.41)
@@
$U_{2}<U<U_{3}:\qquad X(U)>0,\quad X^{\prime\prime}(U)<0\,.$ (IV.42)
The solution is concave, bends back downwards until $U_{3}$. This alternating pattern repeats itself until $U_{m}$. The solution in each of the intervals is monotonic until the (in this interval unique) local minimum/maximum is reached; then it is again monotonic towards the next node. This looks like a distorted HW.
The main reason of the HW-behaviour is that $X(U)$ and $X^{\prime\prime}(U)$ have opposite sign.
Between $U_{1}$ and $U_{m}$ there are exactly $m-1$ such distorted HWs. It is easy to see that this $m-1$ counting is correct by plotting
$m=1\qquad X(U)=-P_{1}(\tanh U)=-\tanh U\,.$ (IV.43)
This solution is monotonic, there is obviously no HW as seen in FIG.#7 of DM-1 . Similarly one can plot
$m=2\qquad X(U)=-\frac{1}{2}+\frac{3}{2}\tanh^{2}U.$ (IV.44)
There is a single HW between $U_{1}=-{\rm arctanh}(1/\sqrt{3})$ and $U_{2}={\rm arctanh}(1/\sqrt{3})$, as seen in FIG.#8 of DM-1  .
@@ IV.5.2 Derived Pöschl-Teller
$\frac{{\rm d}^{2}}{{\rm d}U^{2}}X(U)=-k\tanh U\frac{1}{\cosh^{2}U}X(U)\,.$ (IV.45)
As we have seen, for the HW behaviour it is crucial that $X$ and $X^{\prime\prime}$ have opposite sign. In this case this is satisfied only for $U>0$. However, it turns out that this is not a problem. In our normalization $U(-\infty)=1$ and so the solution starts off positive. But since initially $X^{\prime\prime}$ is also positive, the solution curve is convex, so it bends upwards and does not vanish for negative $U$. After crossing $U=0$ the curve becomes concave and now bends downwards until it reaches the first node $U_{1}$. All nodes of $X^{1}$ are positive ! This means that all our qualitative considerations for the PT case are valid also here: there are exactly $(m-1)$ HWs between $U_{1}$ and $U_{m}$.